1. A sample of gas with an initial volume of 28.4 L at a pressure of 735 mmHg and a temperature of 309 K is compressed to a volume of 14.5 L and warmed to a temperature of 377 K .What is the final pressure of the gas?

2. A cylinder with a moveable piston contains 216 mL of nitrogen gas at a pressure of 1.29 atm and a temperature of 292 K .What must the final volume be for the pressure of the gas to be 1.53 atm atm at a temperature of 337 K ?

3.

All pressure–volume–temperature relationships for gases can be combined into a single relationship known as the combined gas law. This expression can be used when looking at the effect of changes in two of these variables on the third as long as the amount of gas (number of moles) remains constant. To use the combined gas law properly, you must always express the temperatures in kelvins. The combined gas law can be represented as follows:

P1V1T1=P2V2T2

Part 1: A very flexible helium-filled balloon is released from the ground into the air at 20. ∘C. The initial volume of the balloon is 5.00 L, and the pressure is 760. mmHg. The balloon ascends to an altitude of 20 km, where the pressure is 76.0 mmHg and the temperature is −50. ∘C. What is the new volume, V2, of the balloon in liters, assuming it doesn't break or leak?

Part B: Consider 4.00 L of a gas at 365 mmHg and 20. ∘C . If the container is compressed to 2.70 L and the temperature is increased to 38 ∘C , what is the new pressure, P2, inside the container? Assume no change in the amount of gas inside the cylinder.
Answer #1

1)

Given:

Pi = 735 mmHg

Vi = 28.4 L

Vf = 14.5 L

Ti = 309.0 K

Tf = 377.0 K

use:

(Pi*Vi)/(Ti) = (Pf*Vf)/(Tf)

(735 mmHg*28.4 L)/(309.0 K) = (Pf*14.5 L)/(377.0 K)

Pf = 1756 mmHg

= 1756 / 760 atm

= 2.31 atm

Answer: 2.31 atm

Only 1 question at a time please
Answer #2

Appeying idual ga law CiD vnRT R Lusur mau niially P 735 m H4 Vulum R - Coytont 28-4 L Temperature 309 K tinally I4.SL 377k P